I am not quite sure how to even set this up.
Suzi drove home at 60 mph, but her brother Jim, who left at the same time, could drive at only 48 mph. When Suzi arrived, Jim still had 60 miles to go. How far did Suzi drive?

Answer by ankor@dixie-net.com(12701)

Suzi drove home at 60 mph, but her brother Jim, who left at the same time, could drive at only 48 mph. When Suzi arrived, Jim still had 60 miles to go. How far did Suzi drive?
:
Let d = distance driven by Suzi
Then (d-60) = distance driven by jim during the time Suzi completed the trip
:
T = Dist/speed
Suzi's time to complete the trip = Jim's time when he still had 60 mi to go
d/60 = (d-60)/48
:
Clear denominators, mult eq by 240, you have:
4d = 5(d-60)
4d = 5d - 300
4d - 5d = -300
-1d = -300
d = -300/-1
d = + 300 miles
:
Check: At 60 mph, Suzi would take 5 hrs to make that trip of 300 mi.
In the same 5 hrs Jim would travel 5*48 = 240 mi, which leaves 60 mi to complete the same trip.